solve for n; s=n/(n+1)
Found 2 solutions by swincher4391, stanbon:
Answer by swincher4391(1107)

s = n/(n+1)
Multiply both sides by (n+1)
(n+1)*s = n
Solve for 0
(n+1)*s - n = 0
Distribute
sn + s -n = 0
Get all "n" terms on one side, lone s terms on the other.
sn -n = -s
Factor out an n.
n(s-1) = -s
Solve for n.
n = -s/(s-1)

Answer by stanbon(75887)

solve for n; s=n/(n+1)
----
Multiply both sides by n=1 to get:
s(n+1) = n
---
sn + s = n
---
n-sn = s
---
n(1-s) = s
n = s/(1-s)
==================
Cheers,
Stan H.
===================